a heterozygous tomato plant, Rr Wf wf, was test crossed to rrwfwf plant. R =red fruit, r= yellow fruit; and Wf=yellow flowers, wf=white flowers. the following progeny resulted:
77 red fruit, yellow flowers;
73 yellow fruit, white flowers;
28 red fruit, white flowers;
32 yellow fruit, yellow flowers.
a)wat is the recombination frequency btwn the fruit color and flower color loci?
b) what is the linkage phase of the alleles in the dihybrid parent?
Answer #1

20). A). Addition of a pair of electrons from NADPH

Answer:

a).

Hint: Non-recombinant progeny are more common in linkage. Hence, recombinant progeny are red fruit, white flowers and yellow fruit, yellow flowers.

Recombination frequency (RF)= (no. of recombinant progeny / total progeny)100

RF= ( 28+32/210)100 = 28.57%

The recombination frequency btwn the fruit color and flower color loci = 28.57%

b).

Hint: Non-recombinant progeny are more common in linkage.

F1 hybrid = red fruit, yellow flowers / yellow fruit, white flowers = RWf / r wf

The linkage phase is cis (coupling) phase.